OpenStudy (anonymous):

FAN AND METAL!! At Blooming Gardens nursery, they are currently running the following promotion: Buy three trees of the same kind at the regular price and get the fourth at 80% off. Jocelyn is needing some trees for her backyard and decided to buy four red oak trees using the promotion. The regular price for a red oak tree is $390.00. Based on a tax rate of 8.25%, Jocelyn ended up saving $____ by using the promotion.

OpenStudy (anonymous):

tax rate added on trees with promotion : 1248 + .0825(1248) = 1248 + 102.96 = 1350.96 tax rate added on trees without promotion : 1560 + .0825(1560) = 1560 + 128.70 = 1688.70 so the savings with the tax rate included are : 1688.70 - 1350.96 = $ 337.74